CHARACTERIZATION OF STRUCTURAL ISOMERS OF NAPHTHALENE: ELECTRONIC SPEC-TROSCOPY OF Z-PHENYLVINYLACETYLENE

\underline{JOSH J. NEWBY}, CHRISTIAN W. MÜLLER, CHING-PING LIU^{a}, WILLIAM H. JAMES III, EVAN G. BUCHANAN, HSIUPU D. LEE, AND TIMOTHY S. ZWIER, Department of Chemistry, Purdue University, West Lafayette, IN 47907.

In a variety of contexts from planetary atmospheres to combustion processes, the appearance of C_{10}H_{8} is commonly attributed to naphthalene. Often neglected are other stable C_{10}H_{8} isomers that could be formed prior to isomerization to naphthalene, such as phenylvinylacetylene (PVA, 1-phenyl-1-buten-3-yne). The electronic spectroscopy of the E isomer of PVA has already been reported,^{b} but direct isomerization to naphthalene would seem more likely from the Z isomer. Here, we report the jet-cooled laser induced fluorescence and single vibronic level dispersed fluorescence spectra of Z-PVA. The spectra display low frequency vibronic structure reminiscent of that in styrene, attributable to the out-of-plane torsion and bend of the vinylacetylene group. The vibrational assignments are assisted by ab initio calculations and the results are compared with the analogous motions in styrene. Interestingly, the excitation spectrum shows a dramatic decrease in fluorescence intensity about 600 cm^{−1}above the electronic origin. This intensity drop off is attributable to a turn on in nonradiative processes.

aCurrent Address: Institute of Chemistry, Academia Sinica, Taipei, 11529 Taiwan.

bC. P. Liu, J. J. Newby, C. W. Müller, H. D. Lee, and T. S. Zwier J. Phys. Chem. A. 112, 9454 (2008).